OpenStudy (anonymous):

q^2-16/(q+4)^2 simplify by removing factors of 1

OpenStudy (anonymous):

\[\frac{(q+4)(q-4)}{(q+4)^2=\frac{q-4}{q+4}\]

OpenStudy (anonymous):

what the heck is going on with the latex? is it just me?

OpenStudy (saifoo.khan):

(q - 4)(q + 4) ------------ (q + 4)(q+4) q-4 ---- q + 4

OpenStudy (saifoo.khan):

\[\Huge i t-is-workin-for-ME!:D\]